# Controllable Video Generation: A Survey

### v1

@@ Abstract
Index Terms:
Survey, Video Generative Model, Controllable Generation, AIGC
@@ 1 Introduction
As interest in AI-generated content (AIGC) continues to grow, video generation—one of its key domains—has emerged as a prominent focus for both researchers and users alike. Modern video generation methods [1, 2, 3, 4, 5, 6, 7] typically leverage cutting-edge generative paradigms (e.g., diffusion [8, 9] or autoregressive models [10, 11, 12, 13]), combined with large-scale datasets [14, 15, 16], massive model parameters [17, 18, 19], and advanced architectural frameworks [20]. We refer to these models as video generation foundation models, which have significantly advanced the quality of generated videos. The resulting outputs exhibit an unprecedented level of creativity. Despite their impressive generative capabilities, these models often remain constrained by their reliance on text-only conditioning, which limits the degree of control users can exert over the generated content. As a result, users frequently struggle to translate their creative ideas into precise video outputs, thereby diminishing the practical effectiveness of these models in real-world content creation scenarios.
Figure 1: The development trend of controllable video generation methods across seven representative task categories. The line chart illustrates the rapid growth in the number of related works from 2022 to the present, with different categories distinguished by color. Representative works from each period are highlighted above the chart. For instance, VideoCrafter [3] and EchoMimic [4] have achieved 4.9k and 3.9k stars in Github, respectively.
@@
Recent survey papers provide extensively overviews of AI-generated content (AIGC), covering various areas such as video generation based on Generative Adversarial Networks and Variational AutoEncoders [21, 22], diffusion model theories and architectures [23], efficient video diffusion models [24], unified multi-modal video synthesis and understanding [25], video editing [26], foundational video diffusion models [27, 28, 29], and 4D generation applications [30]. While these reviews offer valuable insights, many only provide a cursory examination of video generative models or predominantly concentrate on other modalities. This is a significant gap in the literature regarding controllable video generation. Additionally, existing studies rarely address this topic in various control signals, e.g., depth, sketch, segmentation map, leaving a critical void in understanding the potential for integrating novel conditions into video generative models and their implications for advancing controllable video generation.
@@ 2 Preliminaries
This section first explains the basic theories of generative models. As shown in Fig. 2, we present an illustration of GAN, VAE, Flow-based Models, Diffusion Models, and Autoregressive Models, then we give the taxonomy of the controllable video generation tasks.
2.1 GAN and VAE
Generative Adversarial Networks (GANs) and Variational Autoencoders (VAEs) are two classical models for generative modeling. GAN [31] is based on a minimax game between a generator $G$ and a discriminator $D$. The generator tries to produce realistic samples from random noise vectors $\mathbf{z}\sim p(\mathbf{z})$, while the discriminator attempts to distinguish between real data $\mathbf{x}\sim p(\mathbf{x})$ and generated samples $G(\mathbf{z})$. The objective is given by
$\min_{G}\max_{D}\mathbb{E}_{\mathbf{x}}[\log D(\mathbf{x})]+\mathbb{E}_{\mathbf{z}}[\log(1-D(G(\mathbf{z})))].$
VAE [32] adopts a probabilistic approach, introducing a variational posterior $q(\mathbf{z}|\mathbf{x})$ to approximate the posterior $p(\mathbf{z}|\mathbf{x})$. The training objective is to maximize the Evidence Lower Bound (ELBO) :
$\mathcal{L}(\theta;\mathbf{X})=\mathbb{E}_{q(\mathbf{z}|\mathbf{x};\theta)}[\log p(\mathbf{x}|\mathbf{z};\theta)]-D_{KL}(q(\mathbf{z}|\mathbf{x};\theta)\|p(\mathbf{z})).$
Both models form the foundation for many following developments in generative modeling, including diffusion and flow-based models.
2.2 Diffusion Models
In recent years, diffusion models have emerged as a powerful series of generative models, offering high-quality sample generation. A typical representative is the Denoising Diffusion Probabilistic Models (DDPMs) [8], which add Gaussian noise to pure data and learn to reverse this process at sample generation, following the rule of the Markov Chain.
@@ 2.2 Diffusion Models
    $q(x_{1},\dots,$ $x_{T}|x_{0})=\prod_{t=1}^{T}q(x_{t}|x_{t-1}),$
    $q\left(x_{t}|x_{t-1}\right)$ $=\mathcal{N}\left(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\mathbf{I}\right),$
@@
    $p_{\theta}(x_{0:T})$ $=p({x}_{T})\prod_{t=1}^{T}p_{\theta}({x}_{t-1}|{x}_{t}),$
    $p_{\theta}\left(x_{t-1}|x_{t}\right)$ $=\mathcal{N}\left(x_{t-1};\mu_{\theta}\left(x_{t},t\right),\Sigma_{\theta}\left(x_{t},t\right)\right),$
    where the mean $\mu_{\theta}(x_{t},t)$ and the variance $\Sigma_{\theta}(x_{t},t)$ are parameterized by neural networks trained to predict the denoising transformation from $x_{t}$ to $x_{t-1}$.
To simplify the reverse process, the diffusion models often reparameterize the mean $\mu_{\theta}(x_{t},t)$ using a noise prediction model $\epsilon_{\theta}(x_{t},t)$, which directly estimates the noise added at each timestep instead of recovering the clean data:
@@
2.3 Flow-based Models
Flow-based models are a class of powerful generative models, with the core idea being to learn an accurate, invertible transformation that converts a simple base distribution into a complex target data distribution. This transformation process is typically composed of a series of invertible functions, referred to as “flows”.
Traditional flow-based models [33, 34, 35, 36] rely on discrete sequences of invertible transformations. In contrast, Continuous Normalizing Flows (CNFs) [37] generalize this framework by modeling the transformation as a continuous-time flow governed by an Ordinary Differential Equation (ODE):
$\frac{d}{dt}\phi_{t}(x)$ $=v_{t}\left(\phi_{t}(x)\right),$
$\phi_{0}(x)$ $=x,$
@@ 2.3 Flow-based Models
To improve the efficiency of training CNFs, a modern generative approach called Flow Matching (FM) [9] has been proposed. It offers a new paradigm for learning complex data distribution by directly learning the vector field that transfers samples from the base to the target distribution.
@@
$x_{t}=(1-t)x_{0}+tx_{1},$
and the ground truth velocity at time $t$ is
$v_{t}^{\mathrm{gt}}\left(x_{t}\right)=x_{1}-x_{0},$
@@
This loss encourages the vector field to match the direction and magnitude of the straight-line transport from $x_{0}$ to $x_{1}$ across time.
2.4 Autoregressive Models
Autoregressive (AR) [11, 38, 39, 40, 41, 12, 13] Models generate data by modeling the conditional distribution of each element given the previous ones. Formally, given a sequence $\mathbf{x}=(x_{1},x_{2},\dots,x_{T})$, the joint probability can be factorized as
@@ 2.4 Autoregressive Models
In the early stage of generative models, AR models like PixelCNN [10] generate pixel row by row, to capture local dependencies. While, due to their inherently sequential generation process, suffering from efficiency, they were gradually replaced by diffusion models.
@@
Recently, with the growth of video generation and the strong temporal correlations between frames, AR models have regained attention. Many works adopt hybrid designs, integrating autoregressive priors with diffusion backbones.
2.5 Taxonomy
Controllable video generation is a systematic and complex research area. Most existing studies focus on how to generate videos under specific controls, like pose guidance or subject guidance. According to different control types, this task can be naturally divided into seven sub-tasks, as shown in Fig. 3. To provide a deeper understanding of the mechanisms and offer a comprehensive perspective on this area, we further classify them by their condition types. The key challenge in this field lies in how to inject various conditions into pretrained video generative models. That means these models not only align with text prompts but also cooperate with additional conditions to generate high quality and fidelity videos. In addition, recent research explores multi-condition video generation, where the model is guided by a combination of inputs, such as a reference image, a sparse trajectory, and a motion brush. This setting introduces further complexity, requiring the model to effectively integrate multiple conditions.
@@ 2.5 Taxonomy
Table I: An overview of notable video generative models. Resolution: The frames, width and height of generated video under the default setting. $f$: downsampling factor of autoencoder from pixel space to latent space. : The method uses one-shot data from the Dataset.
@@
Model Venue Param. Resolution f Text Encoder Training Dataset Open Source
@@
[…]
Tune-A-Video [44] ICCV 2023 983M $24\times 512^{2}$ 1, 8x8 CLIP ViT-L/14 DAVIS  [45]
AnimateDiff [46] ICLR 2024 1.4B $16\times 256^{2}$ 1, 8x8 CLIP-ViT-L/14 WebVid-10M [14]
Stable Video Diffusion [47] arXiv 2023 1.5B $14\times 576\times 1024$ 1, 8x8 CLIP-ViT-H/14 Internal Dataset
@@
CogVideoX [48] ICLR 2025 2B&5B $49\times 480\times 720$ 4, 8x8 T5 Internal Dataset
HunyuanVideo [17] arXiv 2024 13B $129\times 720\times 1280$ 4, 8x8 Hunyuan-Large Internal Dataset
StepVideo [18] arXiv 2025 30B $102\times 544\times 992$ 8, 16x16 Hunyuan-CLIP, Step-LLM Internal Dataset
Wan [19] arXiv 2025 1.3B&14B $81\times 720\times 1280$ 4, 8x8 umT5 Internal Dataset
@@
CausVid [49] CVPR 2025 1.4B $120\times 352\times 640$ 4, 8x8 umT5 Internal Dataset
NOVA [50] ICLR 2025 0.3B&0.6B&1.4B $33\times 768\times 480$ 4, 8x8 Phi-2 Internal Dataset
Cosmos [51] arXiv 2025 4B&12B $121\times 704\times 1280$ 4, 8x8 T5-XXL Internal Dataset
UVA [52] RSS 2025 0.5B $16\times 720\times 1280$ 1, 16x16 CLIP-ViT-B/32 Libero10 [53], PushT [54], UMI [55], Human Video [56]
MAGI-1 [57] arXiv 2025 4.5B&24B $24\times 720\times 1280$ 4, 8x8 T5 Internal Dataset

### v2

# These $\equalcontThese authors contributed equally to this work.$

@@ Abstract
keywords:
Video generation, Diffusion model, Controllable generation, Survey
@@ 1 Introduction
As interest in AI-generated content (AIGC) continues to grow, video generation, one of its key domains, has emerged as a prominent focus for both researchers and users alike. Modern video generation methods [ho2022imagen, ma2024followposeposeguidedtexttovideo, chen2023videocrafter1opendiffusionmodels, z1, z2, z3, z4, jiang2025personalized, chen2024echomimiclifelikeaudiodrivenportrait, zhang2025mind, feng2025dit4edit, feng2025follow, geng2024motionprompting, zhu2025aether, zhu2024instantswap, long2025follow] typically leverage cutting-edge generative paradigms (e.g., diffusion [ho2020ddpm, lipman2022flow, wang2025multishotmaster] or autoregressive models [van2016conditional, z5, z6, z7, ramesh2021dalle, chang2022maskgit, z10, tian2024var]), combined with large-scale datasets [Bain21WebVid-10M, chen2024panda-70m, schuhmann2021laion-400m], massive model parameters [kong2025hunyuanvideosystematicframeworklarge, huang2025step, z9, wan2025], and advanced architectural frameworks [Peebles2022DiT]. We refer to these models as video generation foundation models, which have significantly advanced the quality of generated videos. The resulting outputs exhibit an unprecedented level of creativity. Despite their impressive generative capabilities, these models often remain constrained by their reliance on text-only conditioning, which limits the degree of control users can exert over the generated content. As a result, users frequently struggle to translate their creative ideas into precise video outputs, thereby diminishing the practical effectiveness of these models in real-world content creation scenarios.
@@
Figure 1: The development trend of controllable video generation methods across seven representative task categories. The line chart illustrates the rapid growth in the number of related works from 2022 to the present, with different categories distinguished by color. Representative works from each period are highlighted above the chart. For instance, VideoCrafter [chen2023videocrafter1opendiffusionmodels] and EchoMimic [chen2024echomimiclifelikeaudiodrivenportrait] have achieved 4.9k and 3.9k stars in Github, respectively.
Recent survey papers provide extensively overviews of AI-generated content (AIGC), covering various areas such as video generation based on Generative Adversarial Networks and Variational AutoEncoders [10.1145/3487891, yazdani2024comprehensive], diffusion model theories and architectures [xing2024survey], efficient video diffusion models [ulhaq2022efficient], unified multi-modal video synthesis and understanding [zhou2024survey], video editing [sun2024diffusion], foundational video diffusion models [wang2025survey, yang2025videograin, melnik2024video], and 4D generation applications [ren2025gen3c]. While these reviews offer valuable insights, many only provide a cursory examination of video generative models or predominantly concentrate on other modalities. This is a significant gap in the literature regarding controllable video generation. Additionally, existing studies rarely address this topic in various control signals, e.g., depth, sketch, leaving a critical void in understanding the potential for integrating novel conditions into video generative models and their implications for advancing controllable video generation.
@@ 2 Preliminaries
This section first explains the basic theories of generative models. In Fig. 2, we present an illustration of GAN, VAE, Flow-based Models, Diffusion Models, and Autoregressive Models, then we give the taxonomy of the controllable video generation tasks.
GANs and VAEs. Generative Adversarial Networks (GANs) and Variational Autoencoders (VAEs) are two classical models for generative modeling. GAN [goodfellow2014generative] is based on a minimax game between a generator $G$ and a discriminator $D$. The generator tries to produce realistic samples from random noise $\mathbf{z}\sim p(\mathbf{z})$, while the discriminator attempts to distinguish between real data $\mathbf{x}\sim p(\mathbf{x})$ and generated samples $G(\mathbf{z})$. The objective is given by
$\min_{G}\max_{D}$ $\mathbb{E}_{\mathbf{x}}\big[\log D(\mathbf{x})\big]+\mathbb{E}_{\mathbf{z}}\big[\log\big(1-D(G(\mathbf{z}))\big)\big].$
VAE [kingma2013auto] introduces a variational posterior $q(\mathbf{z}|\mathbf{x})$ to approximate the true posterior $p(\mathbf{z}|\mathbf{x})$. The training objective is to maximize the Evidence Lower Bound (ELBO):
$\mathcal{L}(\theta;\mathbf{x})$ $=\mathbb{E}_{q(\mathbf{z}\mid\mathbf{x};\theta)}\big[\log p(\mathbf{x}\mid\mathbf{z};\theta)\big]$
$\quad-D_{\mathrm{KL}}\big(q(\mathbf{z}\mid\mathbf{x};\theta)\,\|\,p(\mathbf{z})\big).$
Both models form the foundation for developments in generative modeling, including diffusion and flow-based models.
Diffusion Models. In recent years, diffusion models have emerged as a powerful series of generative models, offering high-quality sample generation. A typical model is the Denoising Diffusion Probabilistic Model (DDPM) [ho2020ddpm], which add Gaussian noise to pure data and learn to reverse this process at sample generation, following the Markov Chain.
@@
    $q(x_{1},\dots,x_{T}\mid x_{0})$ $=\prod_{t=1}^{T}q(x_{t}\mid x_{t-1}),$
    $q(x_{t}\mid x_{t-1})$ $=\mathcal{N}\!\left(x_{t};\,\sqrt{1-\beta_{t}}\,x_{t-1},\,\beta_{t}\mathbf{I}\right).$
@@
    $p_{\theta}(x_{0:T})$ $=p(x_{T})\prod_{t=1}^{T}p_{\theta}(x_{t-1}\mid x_{t}),$
    $p_{\theta}(x_{t-1}\mid x_{t})$ $=\mathcal{N}\!\left(x_{t-1};\,\mu_{\theta}(x_{t},t),\,\Sigma_{\theta}(x_{t},t)\right).$
where the mean $\mu_{\theta}(x_{t},t)$ and the variance $\Sigma_{\theta}(x_{t},t)$ are parameterized by neural networks trained to predict the denoising transformation from $x_{t}$ to $x_{t-1}$. To simplify the reverse process, the diffusion models often reparameterize the mean $\mu_{\theta}(x_{t},t)$ using a noise prediction model $\epsilon_{\theta}(x_{t},t)$, which directly estimates the noise added at each timestep instead of recovering the clean data:
@@
Flow-based Models. Flow-based models are a class of powerful generative models, with the core idea being to learn an accurate, invertible transformation that converts a simple base distribution into a complex target data distribution. This transformation process is typically composed of a series of invertible functions, referred to as “flows”.
Traditional flow-based models [dinh2014nice, chen2024follow, dinh2016density, kingma2018glow, zhang2025easycontrol, song2022cliptexture, song2023clipvg, song2022clipfont, chen2025transanimate] rely on discrete sequences of invertible transformations. In contrast, Continuous Normalizing Flows (CNFs) [chen2018neural] generalize this framework by modeling the transformation as a continuous-time flow governed by an Ordinary Differential Equation (ODE):
$\frac{d}{dt}\phi_{t}(x)=v_{t}\left(\phi_{t}(x)\right),\quad\phi_{0}(x)=x,$
@@
To improve the efficiency of training CNFs, a modern generative approach called Flow Matching (FM) [lipman2022flow] has been proposed. It offers a new paradigm for learning complex data distribution by directly learning the vector field that transfers samples from the base to the target distribution.
@@
$x_{t}=(1-t)x_{0}+tx_{1},v_{t}^{\mathrm{gt}}\left(x_{t}\right)=x_{1}-x_{0},$
@@
This loss encourages the vector field to match the direction and magnitude of the straight-line transport from $x_{0}$ to $x_{1}$.
Autoregressive Models. Autoregressive (AR) [ramesh2021dalle, Wu2021NÜWAVS, ding2021CogView, yu2022Parti, gafni2022makeascene, chang2022maskgit, tian2024var] Models generate data by modeling the conditional distribution of each element given the previous ones. Formally, given a sequence $\mathbf{x}=(x_{1},x_{2},\dots,x_{T})$, the joint probability can be factorized as
@@
In the early stage, AR models like PixelCNN [van2016conditional] generate pixel row by row, to capture local dependencies. While, due to their inherently sequential generation process, suffering from efficiency, they were gradually replaced by diffusion models. Recently, with the growth of video generation and the strong temporal correlations between frames, AR models have regained attention. Many works adopt hybrid designs, integrating autoregressive priors with diffusion backbones.
3 Method
In this section, we describe the methodology for our literature collection and screening process. We focused on papers published within the past five years (2020–2025), as this period marks the rapid evolution of controllable video generation techniques driven by diffusion models. As illustrated in Fig. 3, we followed the PRISMA [moher2009preferred] guideline with a four-phase procedure to select and review relevant works. To align with our research scope, we established the following three inclusion criteria: C1. Controllable Video Generation: The paper must present methods, frameworks, or systems that enable fine-grained control over video synthesis, beyond standard text-to-video generation. Acceptable control signals include but are not limited to: pose, depth, sketches, bounding boxes, motion trajectories, camera parameters, audio, identity, or multi-modal conditions. C2. Based on Generative Foundation Models: The paper should build upon widely recognized generative paradigms, such as DMs, ARs, GANs, flows, or hybrid architectures. C3. Relevant Contributions: We also include papers that introduce datasets, evaluation benchmarks, systematic analyses, or user-centric tools that facilitate controllable video generation, even if they do not propose new generative models.
Phase 1: Identification. To identify influential works in controllable video generation, we selected a series of publication venues based on their impact in computer vision and graphics: Conference on Computer Vision and Pattern Recognition (CVPR), International Conference on Computer Vision (ICCV), European Conference on Computer Vision (ECCV), Neural Information Processing Systems (NeurIPS), International Conference on Learning Representations (ICLR), International Conference on Machine Learning (ICML), ACM Multimedia (MM), Winter Conference on Applications of Computer Vision (WACV), International Conference on Acoustics, Speech, and Signal Processing (ICASSP), International Conference on Robotics and Automation (ICRA), SIGGRAPH/SIGGRAPH Asia, ACM Transactions on Graphics (TOG), IEEE Transactions on Pattern Analysis and Machine Intelligence (TPAMI), Transactions on Visualization and Computer Graphics (TVCG), Transactions on Multimedia (TMM), and Transactions on Machine Learning Research (TMLR). We also considered high-impact preprints on arXiv to capture the most recent advances. We constructed a set of search queries using keywords reflecting the two core aspects of our survey: (1) controllability and (2) video generation. Example queries used in ACM Digital Library and arXiv search are:
Title/Abstract: ((controllable video) OR (controlled video generation) OR (conditional video synthesis) OR (video control) OR (motion control) OR (camera control) OR (pose-guided video) OR (audio-driven video) OR (text-to-video) OR (image-to-video)) AND (generative model OR diffusion OR VAE OR GAN OR AR).
Figure 3: PRISMA [moher2009preferred] flow diagram of our literature review process. We performed a four-phase screening covering identification, screening, eligibility assessment, and final inclusion, as detailed in Sec. 3. The number of papers retained at each stage is denoted by N.
We included papers published between January 2020 and October 2025, written in English. The initial search yielded n=708 candidate papers.
Phase 2: Screening. We screened the titles and abstracts of the 708 papers against criteria C1–C3. Papers that clearly did not address controllable video generation (e.g., focusing solely on video recognition, compression, or non-generative tasks) were excluded. Additionally, papers that introduced generative models without any mechanism for external control were filtered out. After this phase, 388 papers remained for full-text review. To ensure consistency, two authors independently screened a random subset of 100 papers; the inter-rater agreement measured by Cohen’s Kappa was 0.87.
Phase 3: Eligibility. We assessed the full text of the 388 papers for eligibility. Exclusion at this stage was based on: (a) failure to meet both C1 and C2 upon detailed reading, or (b) limited relevance to the core theme of controllable video generation despite meeting C1 or C2 superficially. For instance, papers that only briefly mentioned controllability as a future direction without substantial technical contribution were excluded. This phase removed 142 papers, leaving 246 papers for final inclusion.
Figure 4: An overall framework for controllable video generation. The seven control modalities form the primary classification axis. The framework connects basic theories (Sec. 2) and video generation foundation models (Sec. 4) with controllable methods and creative supports as key insights (Sec. 5), and maps them to downstream technical tasks and practical applications (Sec. 6).
Phase 4: Taxonomy. To provide a comprehensive and structured overview of the field, we propose a multi-layered framework that captures the entire ecosystem of controllable video generation, spanning from technical foundations to practical applications. As illustrated in Fig. 4, our framework encompasses six interconnected dimensions: Technical Foundations: The generative paradigms that form the basis of modern video synthesis, including GANs, VAEs, DMs, Flows, ARs, and corresponding foundation models. Controllable Modules: The specialized components designed to inject external conditions into pre-trained generative models, such as ControlNet, adapters, encoders, and other modules. Control Modalities: The types of input signals used to guide video generation, categorized into seven distinct classes based on their semantic and structural characteristics. Creative Support: The user-centric capabilities that empower creators through iterative design, real-time preview and adjustment, integrated control, etc. Technical Tasks: The core computer vision tasks enabled by controllable generation, including video completion and inpainting, video composition, video-to-4D generation, etc. Practical Applications: The real-world domains where these technologies demonstrate significant impact, such as film production, autonomous driving, and digital humans.
Within this comprehensive framework, we adopt control modality as our primary taxonomic dimension for three reasons. First, the type of control signal fundamentally determines the architectural design choices, conditioning mechanisms, and technical challenges in video generation pipelines. Second, existing research efforts are predominantly organized around specific control types such as pose-guided animation or audio-driven synthesis, making this classification naturally align with the current literature structure. Third, control modalities serve as the critical bridge between user intent (expressed through various input forms) and the generated visual content. As shown in Fig. 4, we systematically categorize controllable video generation methods into seven modality classes: (1) Structure Control (e.g., pose, depth, sketch, bounding boxes), (2) ID Control (person or subject identity), (3) Image Control (reference images), (4) Temporal Control (flow, trajectory, camera, motion), (5) Audio Control (voice and sound), (6) Other Control (text rendering, style, points, BEV), and (7) Universal Control (multi-condition integration).
The task of controllable generation with text-to-video diffusion models represents a multifaceted and complex field. From the condition perspective, we divide this task into seven sub-stasks . Most works study how to generate video under specific conditions, e.g. pose-guided generation, and depth-guided generation. To reveal the underlying mechanisms and characteristics of these approaches, we further categorize them according to their condition types. The primary challenge in this feild is enabling pretrained T2V diffusion models to handle novel types of conditions and generate videos that are not only aligned with the provided textual prompts but also exhibit high visual quality and strong temporal consistency. Moreover, recent methods inverstagte how to generate videos using multiple conditions, such as given a reference image, sparse trajectory, and motion brush. The main challenge in these tasks is the effective integration of multiple, potentially interacting, conditions, necessitating the capability to express several conditions simultaneously and consistently throughout the generated video sequence.
@@ 3 Method
Table 1: An overview of notable video generative models. Resolution: The frame count, width and height of generated video. $f$: downsampling factor of autoencoder from pixel space to latent space. : The method uses one-shot data from the Dataset.
@@
Model Venue Param. Resolution f Text Encoder Training Dataset Open Source Latency
@@
LVDM [he2022lvdm] arXiv 2022 1.2B $\text{Multi-Frame(16\&1024)}\times 256^{2}$ 4, 8x8 CLIP ViT-L/14 WebVid-2M [Bain21WebVid-10M] 114
Tune-A-Video [wu2023tune] ICCV 2023 983M $24\times 512^{2}$ 1, 8x8 CLIP ViT-L/14 DAVIS  [pont-tuset2017DAVIS] 769
AnimateDiff [guo2023animatediff] ICLR 2024 1.4B $16\times 256^{2}$ 1, 8x8 CLIP-ViT-L/14 WebVid-10M [Bain21WebVid-10M] 122
Stable Video Diffusion [blattmann2023stablevideodiffusionscaling] arXiv 2023 1.5B $14\times 576\times 1024$ 1, 8x8 CLIP-ViT-H/14 Internal Dataset 98
@@
CogVideoX [yang2025cogvideoxtexttovideodiffusionmodels] ICLR 2025 2B&5B $49\times 480\times 720$ 4, 8x8 T5 Internal Dataset 180
HunyuanVideo [kong2025hunyuanvideosystematicframeworklarge] arXiv 2024 13B $129\times 720\times 1280$ 4, 8x8 Hunyuan-Large Internal Dataset 3448
StepVideo [huang2025step] arXiv 2025 30B $204\times 544\times 992$ 8, 16x16 Hunyuan-CLIP, Step-LLM Internal Dataset 743
Wan [wan2025] arXiv 2025 1.3B&14B $81\times 720\times 1280$ 4, 8x8 umT5 Internal Dataset 1864
@@
CausVid [yin2025slowbidirectionalfastautoregressive] CVPR 2025 1.4B $120\times 352\times 640$ 4, 8x8 umT5 Internal Dataset 14
NOVA [deng2025autoregressivevideogenerationvector] ICLR 2025 0.3B&0.6B&1.4B $33\times 768\times 480$ 4, 8x8 Phi-2 Internal Dataset 354
Cosmos [nvidia2025cosmosworldfoundationmodel] arXiv 2025 4B&12B $121\times 704\times 1280$ 4, 8x8 T5-XXL Internal Dataset 119
UVA [li2025unifiedvideoactionmodel] RSS 2025 0.5B $6\times 720\times 1280 1, 1$6x16 CLIP-ViT-B/32 Libero10 [liu2023libero], PushT [chi2024diffusionpolicy], UMI [chi2024UMI], Human Video [clark2025actionfreereasoningpolicygeneralization] ¡1
MAGI-1 [ai2025magi1autoregressivevideogeneration] arXiv 2025 4.5B&24B $24\times 720\times 1280$ 4, 8x8 T5 Internal Dataset 1740
InfinityStar [liu2025infinitystar] arXiv 2025 8B $81\times 720\times 1280$ 4, 8x8 T5 Internal Dataset 58